OpenStudy (anonymous):

solve the inverse variation problem: Y=6 when x=3.find y when x=9

OpenStudy (anonymous):

y=2x x=9 y=2*9=18

OpenStudy (anonymous):

Inverse Variation is \(y = \frac{k}{x}\) First, solve for the constant k with your first values. \[y = \frac kx\]\[6 = \frac k3\]\[k = 18\]Now, use this constant for k and solve for y using the second values.\[y = \frac{18}{9} = ?\]